OpenStudy (anonymous):

what is the circumference of a circle with an radius of 15

jimthompson5910 (jim_thompson5910):

radius is 15 diameter is 30 (multiply radius by 2 to get the diameter)

jimthompson5910 (jim_thompson5910):

Circumference = pi*(diameter) pi is approximately 3.14

OpenStudy (anonymous):

thanks i have another question

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

what is the area of a circle with a circumference of 6 pie

jimthompson5910 (jim_thompson5910):

you are given the circumference

jimthompson5910 (jim_thompson5910):

you need to find the area

jimthompson5910 (jim_thompson5910):

the circumference of a circle can also be found using this formula C = 2*pi*r r is the radius

jimthompson5910 (jim_thompson5910):

if C = 6pi, then what must r be?

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

then you use that value of r and plug it into the area of a circle formula \[\Large A = \pi*r^2\]

OpenStudy (anonymous):

what is the circumference of a circle with an area of 36 pie

jimthompson5910 (jim_thompson5910):

start with the area of a circle formula plug in A = 36pi and solve for r. What do you get?